OpenStudy (anonymous):

What are the lengths of the transverse and conjugate axes of the hyperbola given by the equation x^2/62 - y^2/25=1

OpenStudy (anonymous):

This looks like Conics.

OpenStudy (agent0smith):

Std. form for horizontal hyperbola: \[\frac{ x^2 }{ a^2 } - \frac{ y^2 }{ b^2 }= 1 \] this should help: http://www.tpub.com/math2/Job%202_files/image619.jpg Transverse axis = 2a Conjugate axis = 2b

OpenStudy (anonymous):

it is conics.

OpenStudy (anonymous):

so thetransverse axis would be 62 time 2? an the Conjugat axis 25 times 2? @agent0smith

OpenStudy (agent0smith):

Not quite. \[62 = a^2\]\[25 = b^2\]

OpenStudy (agent0smith):

Make sense @FlyinSolo_424 ?

OpenStudy (anonymous):

wait im confused. Would the transverse just be 62?

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

I figured it out!! The transver I would square root it then times it by two and the 62 was ment to be 64. Therefore, Transvers would be 16 and the conjugate 10. Right ? @agent0smith

OpenStudy (agent0smith):

Correct! Nice job :) And i thought that 62 was meant to be a 64...